Which of the items below would be considered the densest? (Show your step-by-step work for calculating density of each, in the units of g/mL, to support of your final answer)
A piece of wood with a volume of 2.5 L and a mass of 12.5 kg
A ball with a volume of 139 mL and a mass of 93 g
A tire with a volume of 2.12 L and a mass of 4.22 x 104 mg
A wire with a volume of 3.91 x 10-7 L and a mass of 7.93 x 10-1 ng
A block of metal with a volume of 1350 mL and a mass of 1.29 x 103 g
1
Expert's answer
2020-09-14T08:19:17-0400

find the density of all

"p=\frac{m}{V}"

Wood

"p=\frac{12500}{2500}=5"

Ball

"p=\frac{93}{139}=0.67"

Tire

"p=\frac{0.4389}{2120}=0.00021"

Wire

"p=\frac{7.93\times10^{-7}}{3910\times10{-7}}=0.002"

Metal

"p=\frac{132.87}{1350}=0.098"

The wood is the most density